# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

### v2

@@ Introduction
For the $\lambda$ schedule in total, we therefore investigate three variants:
The random $\lambda$ variant samples a different vector from the simplex for each instance, introducing stochastic diversity across training and encouraging robustness.
The Gaussian $\lambda$ variant samples values from a standard normal distribution and normalizes them to lie on the simplex, enabling smooth but non-uniform emphasis over dimensions.
The unimodal $\lambda$ variant uses fixed one-hot vectors to eliminate interpolation and isolate individual preference components.
Together, these ablation studies provide insights into the roles of interpolation, multi-dimensional preference modeling, and the structural benefits of our listwise optimization strategy.
@@ Appendix B B Implementation of _compute_lambda_dpo_loss
This function sequentially performs the following steps:
B.1 Input Tensors and Chunked Log-Probabilities
@@
    Set the chunk size and compute token/sequence log-probabilities:
    To reduce memory usage, we split the batch into chunks of size $C$, denoted by:
[…]
    chunk_size = self.lambda_dpo_chunk_size or input_ids.size(0)
@@
    for start in range(0, input_ids.size(0), chunk_size):
@@
    chunk_ids = input_ids[start:end]
    chunk_mask = attention_mask[start:end]
    chunk_labels = labels[start:end]
@@
    log_probs = F.log_softmax(logits[:, :-1], dim=-1)
    labels_shifted = chunk_labels[:, 1:].clone()
    mask = labels_shifted != self.label_pad_token_id
    labels_shifted[labels_shifted == self.label_pad_token_id] = 0
    label_log_probs = torch.gather(log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
    seq_lp = (label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
    seq_log_probs_list.append(seq_lp)
  3. 3.
    Compute reference log-probabilities (if applicable):
    $\log\pi_{\text{ref}}(y|x)=\begin{cases}\log\text{softmax}(\text{ref\_model}(x)),&\text{if ref model is used}\\ 0,&\text{otherwise}\end{cases}$
@@
    ref_log_probs = F.log_softmax(ref_logits[:, :-1], dim=-1)
    ref_label_log_probs = torch.gather(ref_log_probs, dim=2, index=labels_shifted.unsqueeze(-1)).squeeze(-1)
    ref_lp = (ref_label_log_probs * mask).sum(dim=1) / mask.sum(dim=1)
@@ B.2 Reference Model and Relative Reward
  1. 3.
    Compute the relative reward:
    $r_{\theta}(x,y)=\beta\left(\log\pi_{\theta}(y|x)-\log\pi_{\text{ref}}(y|x)\right)$
    Then reshape:
    $r_{\theta}\in\mathbb{R}^{B\times 4\times 4},\quad\pi^{*}\in\mathbb{R}^{B\times 4\times 4}$
@@
    B = input_ids.size(0) // 16
    r_theta = r_theta.view(B, 4, 4)
    pi_target = pi_target.view(B, 4, 4)
@@ B.3 Listwise Loss by Preference Dimension
    $\ell^{(i)}=-\sum_{j}p^{*(i)}_{j}\cdot\log\left(\frac{\exp(r^{(i)}_{j})}{\sum_{j^{\prime}}\exp(r^{(i)}_{j^{\prime}})}\right)$
@@ B.4 Lambda Weights and Final Loss
  1. 6.
    Sample a random weight vector $\boldsymbol{\lambda}\sim\Delta^{4}$, and compute:
    $\mathcal{L}(\theta)=\sum_{i=1}^{4}\lambda_{i}\ell^{(i)}$
    weights = torch.rand(4, device=input_ids.device)

### v3

@@ Introduction
For the $\lambda$ schedule, we investigate three variants:
[…]
  • Gaussian $\lambda$: we sample each component from a standard normal distribution and then normalize the vector to lie on the simplex, thereby enabling smoothly varying but non-uniform emphasis across dimensions.
  • Unimodal $\lambda$: we use fixed one-hot vectors (e.g., $(1,0,0,\dots)$), effectively disabling interpolation and isolating optimization to a single preference dimension per instance.
Together, these ablations help us understand how balanced weighting, smooth interpolation, and single-objective optimization each contribute to the effectiveness of our listwise $\lambda$-DPO framework.
Comparative Experiment
We conduct a comparative experiment to evaluate the performance of standard DPO and our proposed $\lambda$-DPO under identical training and evaluation setups. Specifically, we focus on two representative models—LLaMA 3.2B Instruct and Qwen 1.8B Chat—and assess generalization across five commonly used benchmarks using the lm-eval-harness framework. Both methods are trained from the same base model (i.e., without supervised fine-tuning), using UltraFeedback preference data.
Main Results
Table 1 reports the accuracy of DPO and $\lambda$-DPO on five representative benchmarks for both LLaMA 3.2 1B Instruct and Qwen 1.8B Chat. On LLaMA, $\lambda$-DPO consistently improves performance across all tasks, with notable gains such as +1.1% on ARC, +1.8% on GSM8K, and +3.6% on TruthfulQA. Qwen 1.8B Chat also sees improvements on most benchmarks, particularly GSM8K (+14.7%) and MMLU (+3.9%). DPO only slightly outperforms $\lambda$-DPO on TruthfulQA in the Qwen setting.
Model Method MMLU ARC HellaSwag TruthfulQA GSM8K
LLaMA 3.2 1B Instruct DPO 0.458 0.364 0.456 0.378 0.696
$\lambda$-DPO 0.460 0.375 0.459 0.414 0.714
Qwen 1.8B Chat DPO 0.379 0.271 0.418 0.366 0.008
$\lambda$-DPO 0.418 0.284 0.428 0.312 0.155
Table 1: Accuracy comparison between DPO and $\lambda$-DPO on five benchmarks using LLaMA 3.2 1B Instruct and Qwen 1.8B Chat. Bold numbers indicate better performance.
Visualization
To further illustrate these results, we provide bar charts comparing DPO and $\lambda$-DPO performance on the two models.
Figure 1: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
Figure 2: Accuracy comparison between DPO and $\lambda$-DPO on Qwen 1.8B Chat.
Performance Analysis and Discussion
As shown in both Table 1 and Figures 14, $\lambda$-DPO consistently improves performance across most tasks. The method yields notable improvements on ARC, GSM8K, and MMLU, while maintaining parity on Winogrande. TruthfulQA is the only benchmark where DPO slightly outperforms $\lambda$-DPO (in the Qwen model).
These results suggest that $\lambda$-DPO effectively captures human preference signals across diverse task types. Its listwise preference modeling contributes to better alignment with human-evaluated quality metrics such as factuality, helpfulness, and fluency.
Nonetheless, for certain adversarial or logic-heavy tasks like TruthfulQA, the benefits of listwise supervision may be less reliable. Integrating $\lambda$-DPO with structured supervision—such as chain-of-thought prompting [23]—could potentially mitigate this gap.
Conclusion of Comparative Analysis
In summary, $\lambda$-DPO demonstrates consistent improvements over DPO in most evaluation scenarios, confirming the effectiveness of listwise preference supervision. Its generalization advantage makes it a strong candidate for broader alignment tasks. Future work may explore hybrid strategies that combine $\lambda$-DPO with structured reasoning methods to further enhance performance on complex benchmarks.
@@ Appendix B B Implementation of _compute_lambda_dpo_loss
This function executes the following procedure:
B.1 Input Tensors and Chunked Computation
@@
    Assume the batch contains 4 preference dimensions and 4 candidate responses (16 sequences in total). Reshape the tensors so that each preference dimension shares the same 4 unique responses:
    num_pref, num_resp = 4, 4
    B = input_ids.size(0) // (num_pref * num_resp)
    seq_len = input_ids.size(1)
    ids_view = input_ids.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
    mask_view = attention_mask.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
    labels_view = labels.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
  3. 3.
    Choose a chunk size (by default equal to ids_view.size(0)) to compute log-probabilities in smaller batches. Optionally compute reference model probabilities:
    chunk_size = self.lambda_dpo_chunk_size or ids_view.size(0)
@@
    for start in range(0, ids_view.size(0), chunk_size):
@@
    chunk_ids = ids_view[start:end]
    chunk_mask = mask_view[start:end]
    chunk_labels = labels_view[start:end]
@@
    seq_lp, valid_len = get_batch_logps(logits, chunk_labels, self.label_pad_token_id)
    seq_log_probs_list.append(seq_lp / valid_len)
@@
    ref_lp, ref_len = get_batch_logps(ref_logits, chunk_labels, self.label_pad_token_id)
    ref_lp = ref_lp / ref_len
@@
  4. 4.
    Concatenate the chunk results and share each set of four log-probabilities across the preference dimensions:
@@
    seq_log_probs = seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
    ref_seq_log_probs = ref_seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
@@ B.2 Reference Model and Relative Reward
  1. 5.
    Compute the relative reward and reshape back to $(B,4,4)$:
@@
    r_theta = r_theta.view(B, num_pref, num_resp)
    pi_target = pi_target.view(B, num_pref, num_resp)
@@ B.4 Lambda Weights and Final Loss
  1. 7.
    The current implementation equally weights the preference dimensions (no random sampling of $\lambda$):
    num_losses = len(listwise_losses)
    weights = torch.ones(num_losses, device=input_ids.device)
@@ C.3 ListwiseDataCollatorWithPadding
Appendix C D Implementation of _compute_ultra_dpo_loss
This routine computes the DPO loss when training on the UltraFeedback dataset, where each example contains four candidate responses with human preference scores. It proceeds as follows:
D.1 Input Aggregation
  1. 1.
    Load the tensors input_ids, attention_mask, labels, and pi_target.
  2. 2.
[…]
  3. 3.
    Apply uniform weights across the four preference dimensions to obtain an aggregated distribution for each example:
    $\text{aggregated}=\sum_{i=1}^{4}\tfrac{1}{4}\,\pi^{*(i)}.$
D.2 Selecting Best and Worst Responses
  1. 4.
    Determine the best response index via $\arg\max$ over aggregated and the worst index via $\arg\min$.
  2. 5.
    Gather the corresponding best and worst sequences from the original batch for further processing.
D.3 Log-Probability Computation
  1. 6.
    Pass the selected sequences through the model, compute token log-probabilities, and average over tokens to obtain sequence log-probabilities.
  2. 7.
    If a reference model is provided, compute its log-probabilities in the same way; otherwise use zero tensors as the reference scores.
D.4 Final DPO Loss
  1. 8.
    Split the policy and reference log-probabilities into best and worst halves.
  2. 9.
    Invoke compute_preference_loss on these values to obtain per-example DPO losses, then average across the batch:
    $\mathcal{L}_{\mathrm{DPO}}=\frac{1}{B}\sum_{b=1}^{B}\mathrm{loss}(y^{\mathrm{best}}_{b},\,y^{\mathrm{worst}}_{b}).$
Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
We provide benchmark accuracy comparisons between DPO and $\lambda$-DPO across multiple language models on the UltraFeedback dataset. Performance is reported across six tasks: mmlu, arc, hellaswag, truthfulqa, wsc273, and gsm8k.
E.1 Qwen-1.8 (Base)
Figure 3 shows that $\lambda$-DPO achieves modest improvements over standard DPO on most tasks using the Qwen-1.8 base model.
Figure 3: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 (base).
E.2 Qwen-1.8 Chat
Figure 4 presents results on Qwen-1.8 Chat, where $\lambda$-DPO consistently improves accuracy, especially on reasoning-heavy benchmarks like arc and hellaswag.
Figure 4: Accuracy comparison between DPO and $\lambda$-DPO on Qwen-1.8 Chat.
E.3 LLaMA 3.2 1B (Base)
As shown in Figure 5, $\lambda$-DPO outperforms DPO slightly on LLaMA 3.2 1B base across four out of six tasks, demonstrating generalization beyond the Qwen series.
Figure 5: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B base model.
E.4 LLaMA 3.2 1B (Instruct)
Figure 6 illustrates results for the Instruct-tuned variant of LLaMA 3.2 1B. The performance gap between $\lambda$-DPO and DPO is generally preserved, indicating robustness to instruction tuning.
Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA 3.2 1B Instruct.
The function thus reduces the four-way UltraFeedback preferences to a single pairwise comparison per prompt and optimizes the model via DPO.